A firm plans to issue $2,000,000 face value of commercial paper for 60 days for short-term cash needs. A dealer will issue the paper at $1,995,000 and will charge an upfront fee of 1% of face value. What is the effective annual rate on this paper?
Answer #1

Effective annual rate = (Face value - net amount realized) /net amount realized

Fee = 2,000,000 *.01 = 20,000

= (2,000,000 - (1,995,000-20,000))/ (1,995,000-20,000)

= 25,000/1,975,000

= 0.0126582278481 * 360/maturity 60 days

= 7.594%